### v2

@@ 3 Deterministic and probabilistic inversion
A partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ is nowhere effectively invertible if
@@ 3.1 Inversion without random oracles
$\mu(\{\hskip 0.85355pt{y\oplus r}:{f(g(y\oplus r))=y}\hskip 0.85355pt\})=\mu(f(2^{\omega}))>0$ (4)
@@
There is a partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ which is
@@
    random-preserving and nowhere effectively invertible
@@
There are partial computable $f,h:\subseteq 2^{\omega}\to 2^{\omega}$ such that
@@
$f:\emptyset^{(\omega)}\mapsto 0^{\omega}\hskip 8.5359pt\textrm{is partial computable.}$.
@@
A nowhere effectively invertible partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ with
positive  $\mathsf{dom}(f)\subseteq\mathsf{ML}$    and    $f(2^{\omega})\cap\mathsf{ML}=\emptyset$
@@
There is a partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ which is
@@
    nowhere effectively invertible
@@
    $\mathsf{dom}(f)\cap\mathsf{ML}=\emptyset$ and $f(2^{\omega})\subseteq\mathsf{ML}$ and $\mu(f(2^{\omega}))>0$.
@@ 3.2 Inversion with random oracles
In computational complexity oneway functions are effectively calculable finite maps that cannot be effectively inverted with positive probability. Levin [17] defined oneway functions $f:\subseteq 2^{\omega}\to 2^{\omega}$ to be partial computable with
@@
$\mu(\mathsf{dom}(f))>0\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\mu\big{(}{\{\hskip 0.85355pt{x\oplus r}:{f(g(f(x),r))=f(x)}\hskip 0.85355pt\}}\big{)}=0$
@@
Let $f:\subseteq 2^{\omega}\to 2^{\omega}$ be partial computable. If
$\mu(\mathsf{dom}(f))>0\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\mu\big{(}{\{\hskip 0.85355pt{x\oplus r}:{f(g(f(x),r))=f(x)}\hskip 0.85355pt\}}\big{)}=0.$
@@
$\mu(f(2^{\omega}))>0\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\mu(\{y\oplus r:f(g(y\oplus r))=y\})=0$
@@
$\mu(f^{-1}(\mathsf{ML}))>0\underset{\not\Leftarrow}{\Rightarrow}\big{(}{\mu(\mathsf{dom}(f))>0\ \wedge\ \mu(f(2^{\omega}))>0}\big{)}.$ (6)
The first implication is trivial and by Proposition 2.5 we may assume that the domain and range of $f$ are $\Pi^{0}_{1}(\emptyset^{\prime})$. Since $\mu(f^{-1}(\mathsf{ML}))>0$ there is a 2-random $x$ such that $f(x)$ is random. By [21, Theorem 4.3] we get that $f(x)$ is a 2-random member of $f(2^{\omega})$ which is $\Pi^{0}_{1}(\emptyset^{\prime})$ so $\mu(f(2^{\omega}))>0$.
@@
If $f:\subseteq 2^{\omega}\to 2^{\omega}$ is partial computable then $\mu(f^{-1}(\mathsf{ML}))>0$ if and only if $f$ extends a random-preserving partial computable function.
@@
$\mu(L)>0\hskip 8.5359pt\textrm{where}\hskip 8.5359ptL:=\{x\oplus r:f(g(f(x),r))=f(x)\}$
@@
$\mu(R)>0\hskip 8.5359pt\textrm{where}\hskip 8.5359ptR:=\{y\oplus r:f(g(y,r))=y\}.$
@@
$\bar{f}(x)=\begin{cases}0f(z)&\text{ if }x=0z\text{ for some }z\\ 1z&\text{ if }x=1(0^{\omega}\oplus z)\text{ for some }z\\ \ \uparrow&\text{ otherwise}\end{cases}$
@@
There is a total computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ such that:
@@
Let $P=\bigcap_{s}P_{s}$ be the $\Pi^{0}_{1}$ class of $\mathsf{PA}$ reals where $P_{s}$ is clopen and $s\mapsto P_{s}$ is computable. Define the total computable surjection $f:2^{\omega}\to 2^{\omega}$ by
@@
Almost all reals $y\in f(2^{\omega})$ are incomplete randoms. For each random $y$ every $w\in f^{-1}(y)$ computes a $\mathsf{PA}$ real. So for almost all $r$ and $y\in f(2^{\omega})$:
@@
for almost all $y\in f(2^{\omega})$. This shows that $f$ is right-oneway. Since
@@
the set $f((2^{\omega}-P)\times 2^{\omega})$ is countable. Since $\mu(P)=0$ by countable additivity there is $y$ with $\mu(f^{-1}(y))>0$. For this $y$ we have
@@
$\mu\left(\{v\oplus r:f(g(f(v),r))=f(v)\}\right)>0.$
@@
4 A left-oneway injection
@@
Let $f:\subseteq 2^{\omega}\to 2^{\omega}$ be a partial computable injection with positive domain. If $x\not\leq_{T}f(x)$ for almost all $x\in\mathsf{dom}(f)$ then $f$ is left-oneway.
@@
This raises the question whether all effective injections are easy to invert.
@@
There is a left-oneway partial computable injection.
@@ 4.1 Outline of the proof
    $f_{s}(\llbracket\sigma\rrbracket):=\bigcup\left\{\llbracket\hat{f}_{s}(\sigma\tau)\rrbracket:\sigma\tau\text{ is an active leaf node of }\hat{f}_{s}\right\}$
@@ 4.2 Construction and verification
  • activation stages $13\langle e,t\rangle+$ 1 where $e$-markers requiring action act
  • expansion stages $23t+$ 2 where we expand $f$ at its active leaf nodes.
@@
  • $j=1-i$, $\hat{f}_{s+1}(\sigma j0^{e+k})=\hat{f}_{s}(\sigma)\rho 00$ and $\sigma j0^{e+k}$ be $e$-satisfied
@@
It remains to show that if $g=\Phi_{e}$ for some $e$ then $g$ inverts $f$ on measure 00.
@@
Transition (a)$\to$(b) is caused by an $e$-marker on some $\sigma$ which will act when transition (b)$\to$(c) occurs. Each such action produces a set of relative measure $2^{-e-k-1}$ where $R_{e}$ is satisfied, so the density of $x_{0}$ in $L$ is $\leq 1-2^{-e-k-1}$, contradicting the hypothesis that $x_{0}$ has density $1$ in $L$.

### v3

@@ 3 Deterministic and probabilistic inversion
A partial computable $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ is nowhere effectively invertible if
@@ 3.1 Inversion without random oracles
$\mu(\{\hskip 0.85355pt{y\oplus r}:{f(g(y\oplus r))=y}\hskip 0.85355pt\})=\mu(f(\{0,1\}^{\omega}))>0$ (4)
@@
There is a partial computable $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ with
@@
    $f$ is random-preserving and nowhere effectively invertible
@@
There are partial computable $f,h:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ with
@@
$f:\emptyset^{(\omega)}\mapsto 0^{\omega}\hskip 8.5359pt\textrm{is partial computable.}$
@@
A nowhere effectively invertible partial computable $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ with
positive  $\mathsf{dom}(f)\subseteq\mathsf{ML}$    and    $f(\{0,1\}^{\omega})\cap\mathsf{ML}=\emptyset$
@@
There is a partial computable $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ with
@@
    $f$ is nowhere effectively invertible
@@
    $\mathsf{dom}(f)\cap\mathsf{ML}=\emptyset$ and $f(\{0,1\}^{\omega})\subseteq\mathsf{ML}$ and $\mu(f(\{0,1\}^{\omega}))>0$.
@@ 3.2 Inversion with random oracles
In computational complexity oneway functions are effectively calculable finite maps that cannot be effectively inverted with positive probability. Levin [17] defined oneway functions $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ to be partial computable with
@@
$\mu(\mathsf{dom}(f))>0\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\mu\big{(}{\{\hskip 0.85355pt{x\oplus r}:{f(g(f(x)\oplus r))=f(x)}\hskip 0.85355pt\}}\big{)}=0$
@@
Let $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ be partial computable. If
$\mu(\mathsf{dom}(f))>0\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\mu\big{(}{\{\hskip 0.85355pt{x\oplus r}:{f(g(f(x)\oplus r))=f(x)}\hskip 0.85355pt\}}\big{)}=0.$
@@
$\mu(f(\{0,1\}^{\omega}))>0\hskip 8.5359pt\textrm{and}\hskip 8.5359pt\mu(\{y\oplus r:f(g(y\oplus r))=y\})=0$
@@
$\mu(f^{-1}(\mathsf{ML}))>0\underset{\not\Leftarrow}{\Rightarrow}\big{(}{\mu(\mathsf{dom}(f))>0\ \wedge\ \mu(f(\{0,1\}^{\omega}))>0}\big{)}.$ (6)
The first implication is trivial and by Proposition 2.5 we may assume that the domain and range of $f$ are $\Pi^{0}_{1}(\emptyset^{\prime})$. Since $\mu(f^{-1}(\mathsf{ML}))>0$ there is a 2-random $x$ such that $f(x)$ is random. By [21, Theorem 4.3] we get that $f(x)$ is a 2-random member of $f(\{0,1\}^{\omega})$ which is $\Pi^{0}_{1}(\emptyset^{\prime})$ so $\mu(f(\{0,1\}^{\omega}))>0$.
@@
If $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ is partial computable then $\mu(f^{-1}(\mathsf{ML}))>0$ if and only if $f$ extends a random-preserving partial computable function.
@@
$\mu(L)>0\hskip 8.5359pt\textrm{where}\hskip 8.5359ptL:=\{x\oplus r:f(g(f(x)\oplus r))=f(x)\}$
@@
$\mu(R)>0\hskip 8.5359pt\textrm{where}\hskip 8.5359ptR:=\{y\oplus r:f(g(y\oplus r))=y\}.$
@@
[…]
@@
There is a total computable $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ such that:
@@
Let $P=\bigcap_{s}P_{s}$ be the $\Pi^{0}_{1}$ class of $\mathsf{PA}$ reals where $P_{s}$ is clopen and $s\mapsto P_{s}$ is computable. Define the total computable surjection $f:\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ by
@@
Almost all reals $y\in f(\{0,1\}^{\omega})$ are incomplete randoms. For each random $y$ every $w\in f^{-1}(y)$ computes a $\mathsf{PA}$ real. So for almost all $r$ and $y\in f(\{0,1\}^{\omega})$:
@@
for almost all $y\in f(\{0,1\}^{\omega})$. This shows that $f$ is right-oneway. Since
@@
the set $f((\{0,1\}^{\omega}-P)\times\{0,1\}^{\omega})$ is countable. Since $\mu(P)=0$ by countable additivity there is $y$ with $\mu(f^{-1}(y))>0$. For this $y$ we have
@@
$\mu\left(\{v\oplus r:f(g(f(v)\oplus r))=f(v)\}\right)>0.$
@@
4 A left-oneway injective function
@@
Let $f:\subseteq\{0,1\}^{\omega}\to\{0,1\}^{\omega}$ be a partial computable injective function with positive domain. If $x\not\leq_{T}f(x)$ for almost all $x\in\mathsf{dom}(f)$ then $f$ is left-oneway.
@@
This raises a question: are all effective injective functions easy to invert?
@@
There is a left-oneway partial computable injective function.
@@ 4.1 Outline of the proof
[…]
@@ 4.2 Construction and verification
  • activation stages $3\langle e,t\rangle+1$ where $e$-markers requiring action act
  • expansion stages $3t+2$ where we expand $f$ at its active leaf nodes.
@@
  • j=1ij=1-i, $\hat{f}_{s+1}(\sigma j0^{e+k})=\hat{f}_{s}(\sigma)\rho 00$ and $\sigma j0^{e+k}$ be $e$-satisfied
@@
It remains to show that if $g=\Phi_{e}$ for some $e$ then $g$ inverts $f$ on measure 0.
@@
[…]